# Spread complexity in saddle-dominated scrambling

### 19 Dec 2023 v1

@@ 1 Introduction
In essence, the exponential growth of OTOCs can manifest even in integrable systems. Thus, relying solely on the exponential growth of OTOCs proves insufficient as a standalone indicator for identifying quantum chaos. Notably, well-established toy models exemplifying this phenomenon (saddle-dominated scrambling) include the Lipkin-Meshkov-Glick (LMG) model Xu:2019lhc and the inverted harmonic oscillator Hashimoto:2020xfr . [^2] See also Bhattacharyya:2020art ; Qu:2021ius ; Qu:2022zwq .
@@
Beyond its potential as a tool for probing quantum chaos, Krylov complexity has spurred extensive research due to its non-trivial connection with the spectral form factor and its implications for holography. [^4] The concept of Krylov complexity is gaining prominence, particularly in the context of holography, as it is well-defined in any quantum theory, overcoming the ambiguities in defining specific elementary gates or establishing tolerances. This interest has led to investigations across various domains, including research in diverse quantum systems Balasubramanian:2022tpr ; Barbon:2019wsy ; Avdoshkin:2019trj ; Dymarsky:2019elm ; Rabinovici:2020ryf ; Cao:2020zls ; Kim:2021okd ; Rabinovici:2021qqt ; Trigueros:2021rwj ; Fan:2022xaa ; Heveling:2022hth ; Bhattacharjee:2022vlt ; Caputa:2022eye ; Caputa:2022yju ; Muck:2022xfc ; Rabinovici:2022beu ; He:2022ryk ; Hornedal:2022pkc ; Alishahiha:2022anw ; Bhattacharjee:2022vlt ; Baek:2022pkt ; Bhattacharjee:2023dik ; Bhattacharyya:2023dhp ; Erdmenger:2023wjg ; Hashimoto:2023swv ; Camargo:2023eev ; Balasubramanian:2023kwd , gauge theories Magan:2020iac , holographic models Jian:2020qpp ; Rabinovici:2023yex , conformal/quantum field theories Dymarsky:2021bjq ; Caputa:2021ori ; Avdoshkin:2022xuw ; Camargo:2022rnt , Lie groups Caputa:2021sib ; Patramanis:2021lkx ; Patramanis:2023cwz ; Chattopadhyay:2023fob , matrix models Iizuka:2023pov ; Iizuka:2023fba , models of quantum quenches Afrasiar:2022efk ; Pal:2023yik and open quantum systems Bhattacharya:2022gbz ; Liu:2022god ; Bhattacharjee:2022lzy ; Bhattacharya:2023zqt .

### 29 Dec 2023 v2

@@ 1 Introduction
In essence, the exponential growth of OTOCs can manifest even in integrable systems. Thus, relying solely on the exponential growth of OTOCs proves insufficient as a standalone indicator for identifying quantum chaos. Notably, well-established toy models exemplifying this phenomenon (saddle-dominated scrambling) include the Lipkin-Meshkov-Glick (LMG) model Xu:2019lhc and the inverted harmonic oscillator Hashimoto:2020xfr . [^2] See also Bhattacharyya:2020art ; Qu:2021ius ; Qu:2022zwq for the circuit complexity and Trunin:2023xmw ; Trunin:2023rwm for the redefined OTOCs.
@@
Beyond its potential as a tool for probing quantum chaos, Krylov complexity has spurred extensive research due to its non-trivial connection with the spectral form factor and its implications for holography. [^4] The concept of Krylov complexity is gaining prominence, particularly in the context of holography, as it is well-defined in any quantum theory, overcoming the ambiguities in defining specific elementary gates or establishing tolerances. This interest has led to investigations across various domains, including research in diverse quantum systems Balasubramanian:2022tpr ; Barbon:2019wsy ; Avdoshkin:2019trj ; Dymarsky:2019elm ; Rabinovici:2020ryf ; Cao:2020zls ; Kim:2021okd ; Rabinovici:2021qqt ; Trigueros:2021rwj ; Fan:2022xaa ; Heveling:2022hth ; Bhattacharjee:2022vlt ; Caputa:2022eye ; Caputa:2022yju ; Muck:2022xfc ; Rabinovici:2022beu ; He:2022ryk ; Hornedal:2022pkc ; Alishahiha:2022anw ; Bhattacharjee:2022vlt ; Baek:2022pkt ; Bhattacharjee:2023dik ; Bhattacharyya:2023dhp ; Erdmenger:2023wjg ; Hashimoto:2023swv ; Camargo:2023eev ; Balasubramanian:2023kwd ; Bhattacharya:2023yec ; Bhattacharjee:2022qjw , gauge theories Magan:2020iac , holographic models Jian:2020qpp ; Rabinovici:2023yex , conformal/quantum field theories Dymarsky:2021bjq ; Caputa:2021ori ; Avdoshkin:2022xuw ; Camargo:2022rnt ; Adhikari:2022whf ; Vasli:2023syq , Lie groups Caputa:2021sib ; Patramanis:2021lkx ; Patramanis:2023cwz ; Chattopadhyay:2023fob , matrix models Iizuka:2023pov ; Iizuka:2023fba , models of quantum quenches Afrasiar:2022efk ; Pal:2023yik and open quantum systems Bhattacharya:2022gbz ; Liu:2022god ; Bhattacharjee:2022lzy ; Bhattacharya:2023zqt ; Bhattacharjee:2023uwx .